# Creative Preference Optimization

### v1

@@ Abstract
Mete Ismayilzada1,2, Antonio Laverghetta Jr.3, Simone A. Luchini3, Reet Patel3, Antoine Bosselut1, Lonneke van der Plas 2 Roger Beaty 3 1EPFL, 2Università della Svizzera Italiana, 3Pennsylvania State University mahammad.ismayilzada@epfl.ch
@@ 1 Introduction
Large Language Models (LLMs) have made significant progress across a broad range of natural language generation tasks Team et al. (2023); Zhao et al. (2025); Bubeck et al. (2023); Wei et al. (2022); Brown et al. (2020). However, whether LLMs exhibit true human-like creativity i.e the ability to produce novel (i.e., original), high-quality (i.e. useful) and surprising (i.e. unexpected) ideas Simonton (2012); Boden (2004) remains unclear. Research on the creativity of LLMs has found mixed results, with some reporting that LLMs are more creative than humans Bellemare-Pepin et al. (2024); Zhao et al. (2024), others reporting that they are less creative Koivisto and Grassini (2023); Chakrabarty et al. (2024); Ismayilzada et al. (2024b), and some finding their creativity to be on par with each other Stevenson et al. (2022); Góes et al. (2023); Gilhooly (2024). However, past research has also found that the high LLM performance can be attributed to the artificial nature of the creativity tasks Ismayilzada et al. (2024a) commonly employed to evaluate LLMs such as the Alternative Uses Task Guilford (1967) or to the remarkable creativity of human-written texts on the web Lu et al. (2024). Consequently, LLMs have been shown to often lack novelty and surprise in their generations Ismayilzada et al. (2024a, b); Zhang et al. (2025); Tian et al. (2024); Chakrabarty et al. (2024) and produce significantly less diverse content compared to humans Padmakumar and He (2023); Anderson et al. (2024); Kirk et al. (2023); Xu et al. (2024); O’Mahony et al. (2024); Zhang et al. (2024); Wenger and Kenett (2025). These tendencies limit the utility of LLMs for creative tasks, such as story generation and creative problem solving that often require longer responses and “out-of-the-box” thinking Tian et al. (2023); Huang et al. (2024); Chen et al. (2024).
@@
Recent research has proposed some methods for improving creativity of LLMs, often targeting diversity aspect alone Wong et al. (2024); Hayati et al. (2023); Chung et al. (2023); Franceschelli and Musolesi (2024); Zhang et al. (2024); Wang et al. (2024b); Zhou et al. (2025); Lanchantin et al. (2025); Chung et al. (2025) or focusing on a single creativity task Tian et al. (2023); Nair et al. (2024); Summers-Stay et al. (2023). However, creativity is a multifaceted ability that also encompasses novelty, surprise, and quality and manifests itself in a wide range of tasks. Consequently, it has been argued that methods promoting creativity improvements should consider multiple dimensions of creativity together across several creative tasks Ismayilzada et al. (2024a). Hence, the broader challenge of enhancing overall creativity in LLM outputs largely remain underexplored.
@@
To this end, we propose a novel approach to directly optimize for creativity in language model generation through preference learning Ouyang et al. (2022); Rafailov et al. (2023). Recent works targeting improvement in LLM creativity have mainly focused on black-box techniques to elicit creative outputs through input-level (e.g., prompting) Tian et al. (2023); Mehrotra et al. (2024); Nair et al. (2024); Summers-Stay et al. (2023) and output-level strategies (e.g., creative decoding) Franceschelli and Musolesi (2024); Meister et al. (2023). However, these methods are inherently limited to the fixed creative capacity of language models and are not designed to optimize for fine-grained dimensions of creativity. Recently, motivated by the negative impact of the preference alignment techniques on the diversity of LLM outputs Padmakumar and He (2023); Anderson et al. (2024); Kirk et al. (2023); O’Mahony et al. (2024); West and Potts (2025), few works have suggested directly modifying the preference optimization methods to promote output diversity Lanchantin et al. (2025); Chung et al. (2025). Inspired by these approaches, we design a new optimization strategy that injects signals from multiple dimensions of creativity into the preference modeling objective in a modular fashion. Specifically, we integrate the novelty, diversity, surprise and quality dimensions of creativity into the training objective of direct preference optimization (DPO) Rafailov et al. (2023), with weighted composition that allow balancing each dimension’s contribution. We call this method creative preference optimization (CrPO) and provide its conceptual illustration in Figure 1 with full details in Section 3.
@@
We test the efficacy of CrPO using MuCE (Multitask Creativity Evaluation), our newly curated large-scale dataset of prompt-response pairs annotated with human preferences across a diverse range of creative tasks in multiple languages. While previous work has largely evaluated creativity improvements on a narrow range of tasks like story generation Chung et al. (2025); Lanchantin et al. (2025) or creative problem solving Tian et al. (2023), MuCE enables us to test whether our methods truly generalize across a diverse range of creativity assessments. Our results show that Llama-3.1-8B-Instruct AI@Meta (2024) and Mistral-7B-Instruct-v0.3 Jiang et al. (2023) trained using CrPO outperform the same models trained using only supervised fine-tuning (SFT) or DPO without any creativity injections, as well as existing LLMs such as GPT-4o, generating more novel, diverse, and surprising outputs than all the baselines while maintaining high quality.
@@ 3 Creative Preference Optimization
According to its three-criterion definition, creativity involves the generation of novel, high-quality, and surprising ideas Simonton (2012); Boden (2004); Runco and Jaeger (2012). Moreover, creative outputs tend to be highly diverse across individuals Anderson et al. (2024). Therefore, to promote overall creativity in LLM outputs, we propose to inject unsupervised metrics related to each dimension of creativity into the loss functions of standard preference optimization methods. We use direct preference optimization (DPO) Rafailov et al. (2023) to illustrate our modifications to the loss function. Recall that in the standard formulation of DPO, a policy model ($p_{\theta}$) is directly optimized on a dataset of $(x,y^{w},y^{l})$ where $x$, $y^{w}$ and $y^{l}$ refer to the model input (i.e. prompt), preferred (i.e. chosen) model response and dispreferred (i.e. rejected) model response, respectively. Using the ratio between the policy model’s likelihood and that of the reference SFT model ($p_{SFT}$) as an implicit reward, the training objective of DPO is defined as follows:
@@
A challenge with standard preference optimization methods is that they may significantly reduce the diversity of the responses LLMs generate, as the loss function encourages models to generate preferred responses even if they are not very creative West and Potts (2025); Padmakumar and He (2023); Anderson et al. (2024); Kirk et al. (2023); Xu et al. (2024); O’Mahony et al. (2024); Zhang et al. (2024); Wenger and Kenett (2025). Existing approaches to address this in the preference optimization objective have centered around curating a preference data based on various diversity metrics Lanchantin et al. (2025) or incorporating extra regularization terms that encourage diverse generations while balancing quality Chung et al. (2025). For example, the recently proposed Diversified DPO (DDPO) method adds a scalar diversity term $\delta^{w}$ (i.e. diversity score of the preferred response) into the DPO loss Chung et al. (2025):
@@ 3.1 Diversity
We use an inverse homogenization metric from Padmakumar and He (2023) similar to Chung et al. (2025). Specifically, given a prompt $x$ and a set of (preferred) responses for $x$ denoted as $Y_{x}$, we compute the diversity score of any particular preferred response as the average pairwise semantic distance to all the other preferred responses in $Y_{x}$:
@@ 3.2 Novelty
We use a novelty metric similar to Karampiperis et al. (2014) where the novelty of a text is defined as the absolute difference between the average pairwise semantic distances of words in the text and those of a reference corpus of texts. In particular, we define the set of preferred responses to a prompt $x$ as a reference corpus ($Y_{x}$) and define the novelty of a preferred response as follows:
@@
Here $T$ refers to a piece of text, $T_{i}$ to the word $i$ in the set of unique words in $T$ denoted as $|T|$ and $DSI(\cdot)$ is divergent semantic integration, the average pairwise semantic distances of words in a text Johnson et al. (2022).
@@ 3.3 Surprise
We use Shannon surprise – the negative log-likelihood of the text — which has been widely used as a measure of surprise in prior work Bunescu and Uduehi (2022); Modirshanechi et al. (2022); Kuznetsova et al. (2013). More specifically, given a prompt $x$, we define the surprise of a particular response as the exponentiated negative log-likelihood of the response (i.e. perplexity) conditioned on the prompt $x$ and under some reference model $S$ as follows:
@@ 3.4 Quality
Although a general quality scoring method is hard to define, reward models that are trained to output a high score to preferred answers can be used as a proxy Zhang et al. (2025); Lambert et al. (2024). In particular, we define the quality of a preferred response given a prompt $x$ as the score assigned by some reward model $R$: $\gamma^{w}=R(y^{w}|x)$.
@@ 4 The MuCE Dataset
Every response in MuCE was rated for creativity by at least two raters, and in some cases up to 75 employing a missing-raters design Forthmann et al. (2025). While it is common practice to measure creativity using multiple independent raters, individual raters may deliver unhelpful or noisy ratings if they did not understand the task instructions, had a different understanding of the rating criteria, or for other reasons Forthmann et al. (2017). To account for this, we followed best practices for subjective scoring tasks by employing Judge Response Theory Myszkowski and Storme (2019) to check for raters whose ratings were uninformative in an information-theoretic sense. We fit JRT models to each task within MuCE, which gave us an information function for each rater across tasks. We then input the results from the JRT into a genetic algorithm Schroeders et al. (2016) which identified a subset of raters per dataset that maximized the per dataset rater information function. [^2] While ensuring that the algorithm kept at least two raters per dataset. This process dropped uninformative raters from each dataset, enhancing the quality of the final creativity ratings. The individual rater’s scores were aggregated via factor scores, as is best practice in creativity assessment Silvia (2011), and we rescaled the factor-transformed creativity scores into the integer range 10-50 as is done for prior work in automated creativity assessment Organisciak et al. (2023). From this dataset, we create multiple data splits for training and testing. Full details about the dataset construction are in Appendix A.
@@ 5.1 SFT and Preference Datasets
While our MuCE dataset contains samples for multiple languages, we focus on showing the effectiveness of CrPO on the English subset in this work and leave experiments using the full dataset as future work. From the base English MuCE dataset, we generate a preference dataset by creating tuples of preferred and rejected responses to the same prompt, treating the response that received the higher creativity score as the preferred one. Past work has shown that data quality is one of the main factors behind preference model performance Liu et al. (2024); Deng et al. (2025); Wang et al. (2024a). Therefore, we curate a high-quality SFT dataset of $5,275$ samples (MuCE-SFT) and preference dataset of $42,058$ samples (MuCE-Pref) from the base MuCE which we detail in Appendix B.
@@ 5.3 Evaluation
We evaluate all models across several dimensions of creativity on held-out prompts of various tasks and two held-out tasks. More specifically, we use $6$ held-out prompts from Real-Life Creative Problem Solving, Alternate Uses of Objects, Design Solutions, Hypothesis Generation, and Metaphors tasks, and $9$ prompts from two held-out tasks of Poems and Sentence Completion. For each prompt, we generate $16$ responses from each model by varying the $temperature$, $top_{p}$, and $top_{k}$ decoding parameters. Our final held-out evaluation suite contains $224$ samples. We evaluate the responses on the dimensions of novelty, diversity, and surprise using the metrics described in Section 3. Additionally, to study the tradeoff between creativity and quality, we train a reward model on our preference dataset using instruction tuned Gemma-2-9b Google (2024a) and use it to score the overall quality of model generations. More details about the evaluation setup can be found in Appendix D.
@@
As baselines, we use the base models Llama-3.1-8B-Instruct AI@Meta (2024) and Mistral-7B-Instruct-v0.3 Jiang et al. (2023), SFT models which are the base models supervised fine-tuned on MuCE-SFT, vanilla DPO model trained on top of the SFT model using the MuCE-Pref dataset without any creativity injections and three closed-source instruction-tuned LLMs, namely GPT-4o OpenAI (2024), Claude-3.7-Sonnet Anthropic (2025), and Gemini-2.0-Flash Google (2024b).

### v2

@@ Abstract
Mete Ismayilzada1,2, Antonio Laverghetta Jr.*3, Simone A. Luchini4, Reet Patel4, Antoine Bosselut1, Lonneke van der Plas †2, Roger E. Beaty †4 1EPFL, 2Università della Svizzera Italiana, 3Wesleyan University, 4Pennsylvania State University mahammad.ismayilzada@epfl.ch
[…]
@@ 1 Introduction
Large Language Models (LLMs) have made significant progress across a broad range of natural language generation tasks Team et al. (2023); Zhao et al. (2025); Bubeck et al. (2023); Wei et al. (2022); Brown et al. (2020). However, whether LLMs exhibit true human-like creativity, i.e, the ability to produce novel (i.e., original), high-quality (i.e., useful) and surprising (i.e, unexpected) ideas Simonton (2012); Boden (2004) remains unclear. Research on the creativity of LLMs has found mixed results, with some reporting that LLMs are more creative than humans Bellemare-Pepin et al. (2024); Zhao et al. (2024); Stevenson et al. (2022), others reporting that they are less creative Koivisto and Grassini (2023); Chakrabarty et al. (2024); Ismayilzada et al. (2024b), and some finding their creativity to be on par with each other Góes et al. (2023); Gilhooly (2024). However, past research has also found that the high LLM performance can be attributed to the artificial nature of the creativity tasks Ismayilzada et al. (2024a) commonly employed to evaluate LLMs, such as the Alternative Uses Task Guilford (1967) or to the remarkable creativity of human-written texts on the web Lu et al. (2024). Consequently, LLMs have been shown to often lack novelty and surprise in their generations Ismayilzada et al. (2024a, b); Zhang et al. (2025); Tian et al. (2024); Chakrabarty et al. (2024) and produce significantly less diverse content compared to humans Padmakumar and He (2023); Anderson et al. (2024); Kirk et al. (2023); Xu et al. (2024); O’Mahony et al. (2024); Zhang et al. (2024); Wenger and Kenett (2025). These tendencies limit the utility of LLMs for creative tasks, such as story generation and creative problem solving that often require longer responses and “out-of-the-box” thinking Tian et al. (2023); Huang et al. (2024); Chen et al. (2024).
@@
Recent research has proposed some methods for improving the creativity of LLMs, often targeting the diversity aspect alone Wong et al. (2024); Hayati et al. (2023); Chung et al. (2023); Franceschelli and Musolesi (2024); Zhang et al. (2024); Wang et al. (2024b); Zhou et al. (2025); Lanchantin et al. (2025); Chung et al. (2025) or focusing on a single creativity task Tian et al. (2023); Nair et al. (2024); Summers-Stay et al. (2023). However, creativity is a multifaceted ability that also encompasses novelty, surprise, and quality and manifests itself in a wide range of tasks. Consequently, it has been argued that methods promoting creativity improvements should consider multiple dimensions of creativity together across several creative tasks Ismayilzada et al. (2024a). Hence, the broader challenge of enhancing overall creativity in LLM outputs largely remains underexplored.
@@
To this end, we propose a novel approach to directly optimize for creativity in language model generation through preference learning Ouyang et al. (2022); Rafailov et al. (2023). Recent works targeting improvement in LLM creativity have mainly focused on black-box techniques to elicit creative outputs through input-level (e.g., prompting) Tian et al. (2023); Mehrotra et al. (2024); Nair et al. (2024); Summers-Stay et al. (2023) and output-level strategies (e.g., creative decoding) Nguyen et al. (2024); Franceschelli and Musolesi (2024); Meister et al. (2023). However, these methods are inherently limited to the fixed creative capacity of language models and are not designed to optimize for fine-grained dimensions of creativity. Recently, motivated by the negative impact of the preference alignment techniques on the diversity of LLM outputs Padmakumar and He (2023); Anderson et al. (2024); Kirk et al. (2023); O’Mahony et al. (2024); West and Potts (2025), few works have suggested directly modifying the preference optimization methods to promote output diversity Lanchantin et al. (2025); Chung et al. (2025). Inspired by these approaches, we design a new optimization strategy that injects signals from multiple dimensions of creativity into the preference modeling objective in a modular fashion. Specifically, we integrate the novelty, diversity, surprise, and quality dimensions of creativity into the training objective of direct preference optimization (DPO) Rafailov et al. (2023), with weighted composition that allows balancing each dimension’s contribution. We call this method creative preference optimization (CrPO) and provide its conceptual illustration in Figure 1 with full details in Section 3.
@@
We test the efficacy of CrPO using MuCE (Multitask Creativity Evaluation), our newly curated large-scale dataset of prompt-response pairs annotated with human preferences across a diverse range of creative tasks in multiple languages. While previous work has largely evaluated creativity improvements on a narrow range of tasks like story generation Ismayilzada et al. (2024b); Chung et al. (2025); Lanchantin et al. (2025) or creative problem solving Tian et al. (2023), MuCE enables us to test whether our methods truly generalize across a diverse range of creativity assessments. Our results show that Llama-3.1-8B-Instruct AI@Meta (2024) and Mistral-7B-Instruct-v0.3 Jiang et al. (2023) trained using CrPO outperform the same models trained using only supervised fine-tuning (SFT) or DPO without any creativity injections, as well as existing LLMs such as GPT-4o, generating more novel, diverse, and surprising outputs than all the baselines while maintaining high quality. We publicly release our code, models, and data for future research [^1] https://www.mete.is/creative-preference-optimization/.
@@ 3 Creative Preference Optimization
According to its three-criterion definition, creativity involves the generation of novel, high-quality, and surprising ideas Simonton (2012); Boden (2004); Runco and Jaeger (2012). Moreover, creative outputs tend to be highly diverse across individuals Anderson et al. (2024). Therefore, to promote overall creativity in LLM outputs, we propose to inject unsupervised metrics related to each dimension of creativity into the loss functions of standard preference optimization methods. We use direct preference optimization (DPO) Rafailov et al. (2023) to illustrate our modifications to the loss function. Recall that in the standard formulation of DPO, a policy model ($p_{\theta}$) is directly optimized on a dataset of $(x,y^{w},y^{l})$ where $x$, $y^{w}$ and $y^{l}$ refer to the model input (i.e. prompt), preferred (i.e. chosen) model response and dispreferred (i.e. rejected) model response, respectively. Using the ratio between the policy model’s likelihood and that of the reference SFT model ($p_{SFT}$) as an implicit reward, the training objective of DPO is defined as follows Rafailov et al. (2023):
@@
A challenge with standard preference optimization methods is that they may significantly reduce the diversity of the responses LLMs generate, as the loss function encourages models to generate preferred responses even if they are not very creative West and Potts (2025); Padmakumar and He (2023); Anderson et al. (2024); Kirk et al. (2023); Xu et al. (2024); O’Mahony et al. (2024); Zhang et al. (2024); Wenger and Kenett (2025). Existing approaches to address this in the preference optimization objective have centered around curating preference data based on various diversity metrics Lanchantin et al. (2025) or incorporating extra regularization terms that encourage diverse generations while balancing quality Chung et al. (2025). For example, the recently proposed Diversified DPO (DDPO) method adds a scalar diversity term $\delta^{w}$ (i.e. diversity score of the preferred response) into the DPO loss Chung et al. (2025):
@@ 3.1 Diversity
Diversity is defined as the pairwise differences between artifacts, and semantic distance is typically used to measure the difference Chung et al. (2025); Ismayilzada et al. (2024b); Padmakumar and He (2023). We use an inverse homogenization metric from Padmakumar and He (2023) similar to Chung et al. (2025). Specifically, given a prompt $x$ and a set of (preferred) responses for $x$ denoted as $Y_{x}$, we compute the diversity score of any particular preferred response as the average pairwise semantic distance to all the other preferred responses in $Y_{x}$:
@@ 3.2 Novelty
Novelty is typically defined as the measure of how different an artifact is from other known artifacts in its class Maher (2010), and semantic distance-based metrics have been established as a good proxy in creativity research Johnson et al. (2022); Beaty and Johnson (2021); Dunbar and Forster (2009); Harbinson and Haarman (2014); Karampiperis et al. (2014). We use a novelty metric similar to Karampiperis et al. (2014) where the novelty of a text is defined as the absolute difference between the average pairwise semantic distances of words in the text and those of a reference corpus of texts. In particular, we define the set of preferred responses to a prompt $x$ as the reference corpus ($Y_{x}$) and define the novelty of a preferred response as follows:
@@
Here $T$ refers to a piece of text, $T_{i}$ to the word $i$ in the set of unique words in $T$ denoted as $|T|$, and $DSI(\cdot)$ is divergent semantic integration, the average pairwise semantic distances of words in a text Johnson et al. (2022).
@@ 3.3 Surprise
Surprise or unexpectedness has many definitions in the cognitive science literature Modirshanechi et al. (2022), but it is generally characterized by deviation from the expected. We use Shannon surprise – the negative log-likelihood of the text — which has been widely used as a measure of surprise in prior work Bunescu and Uduehi (2022); Modirshanechi et al. (2022); Kuznetsova et al. (2013). More specifically, given a prompt $x$, we define the surprise of a particular response as the exponentiated negative log-likelihood of the response (i.e. perplexity) conditioned on the prompt $x$ and under some reference model $S$ as follows:
@@ 3.4 Quality
Although a general quality scoring method is hard to define as it is highly domain-dependent, reward models that are trained to output a high score to preferred answers are increasingly being used to assess the overall quality of language model outputs and align them to human preference Zhang et al. (2025); Lambert et al. (2024). In particular, we define the quality of a preferred response given a prompt $x$ as the score assigned by some reward model $R$: $\gamma^{w}=R(y^{w}|x)$.
@@ 4 The MuCE Dataset
Every response in MuCE was rated for creativity by at least two raters, and in some cases up to 75 employing a missing-raters design Forthmann et al. (2025). While it is common practice to measure creativity using multiple independent raters, individual raters may deliver unhelpful or noisy ratings if they did not understand the task instructions, had a different understanding of the rating criteria, or for other reasons Forthmann et al. (2017). To account for this, we followed best practices for subjective scoring tasks by employing Judge Response Theory Myszkowski and Storme (2019) to check for raters whose ratings were uninformative in an information-theoretic sense. We fit JRT models to each task within MuCE, which gave us an information function for each rater across tasks. We then input the results from the JRT into a genetic algorithm Schroeders et al. (2016) which identified a subset of raters per dataset that maximized the per dataset rater information function. [^3] While ensuring that the algorithm kept at least two raters per dataset. This process dropped uninformative raters from each dataset, enhancing the quality of the final creativity ratings. The individual rater’s scores were aggregated via factor scores, as is best practice in creativity assessment Silvia (2011), and we rescaled the factor-transformed creativity scores into the integer range 10-50 as is done for prior work in automated creativity assessment Organisciak et al. (2023). Full details about the dataset construction are in Appendix A.
@@ 5.1 SFT and Preference Datasets
While our MuCE dataset contains samples for multiple languages, we focus on showing the effectiveness of CrPO on the English subset in this work and leave experiments using the full dataset as future work. From the base English MuCE dataset, we generate a preference dataset by creating tuples of preferred and rejected responses to the same prompt, treating the response that received the higher creativity score as the preferred one. Past work has shown that data quality is one of the main factors behind preference model performance Liu et al. (2024); Deng et al. (2025); Wang et al. (2024a). Therefore, we curate a high-quality SFT dataset of $5,285$ samples (MuCE-SFT) and preference dataset of $42,058$ samples (MuCE-Pref) from the base MuCE which we detail in Appendix B.
@@ 5.3 Evaluation
We evaluate all models across several dimensions of creativity on held-out prompts of various tasks and two held-out tasks. More specifically, we use $5$ held-out prompts from Real-Life Creative Problem Solving, Alternate Uses of Objects, Design Solutions, Hypothesis Generation, and Metaphors tasks, and $9$ prompts from two held-out tasks of Poems and Sentence Completion. For each prompt, we generate $16$ responses from each model by varying the temperaturetemperature, $top_{p}$, and $top_{k}$ decoding parameters. Our final held-out evaluation suite contains $224$ samples. We evaluate the responses on the dimensions of novelty, diversity, and surprise using the metrics described in Section 3. Additionally, to study the tradeoff between creativity and quality, we train a reward model on our preference dataset using instruction tuned Gemma-2-9b Google (2024a) and use it to score the overall quality of model generations. More details about the evaluation setup can be found in Appendix D.
@@
As baselines  [^5] We note that DDPO Chung et al. (2025) is a special case of our method, and the model trained with it is equivalent to our CrPO-div variant. For this reason, we do not treat it as a separate baseline. Moreover, since DDPO has already been shown to outperform the Diverse Preference Optimization method Lanchantin et al. (2025), a direct comparison is unnecessary by transitivity., we use the base models Llama-3.1-8B-Instruct AI@Meta (2024) and Mistral-7B-Instruct-v0.3 Jiang et al. (2023), SFT models which are the base models supervised fine-tuned on MuCE-SFT, a vanilla DPO model trained on top of the SFT model using the MuCE-Pref dataset without any creativity injections and three closed-source instruction-tuned LLMs, namely GPT-4o OpenAI (2024), Claude-3.7-Sonnet Anthropic (2025), and Gemini-2.0-Flash Google (2024b). We also consider additional baselines such as "Brainstorm, then select", a creative prompting approach that has shown improvement on creativity scores in the Alternative Uses of Objects Task (AUT) Summers-Stay et al. (2023) and min-p sampling, a decoding strategy that directly promotes output creativity Nguyen et al. (2024). Evaluation details for these baselines can be found in Appendix D.3.